OpenStudy (anonymous):

solve 2x^2 + 8x + 20 = 0 by completing the square

OpenStudy (solomonzelman):

2x^2 + 8x + 20 = 0 factor out of 2, 2(x^2 + 4x + 10) = 0 divide both sides by 2, x^2 + 4x + 10 = 0 subtract 6 from both sides, -6 -6 x^2+4x+4=-6 complete the square now, can you go on?